# Short Paper

### v1

@@ Abstract
In the landscape of generative artificial intelligence, diffusion-based models present challenges for socio-technical systems in data requirements and privacy. Traditional approaches like federated learning distribute the learning process but strain individual clients, especially with constrained resources (e.g., edge devices). In response to these challenges, we introduce CollaFuse, a novel framework inspired by split learning. Tailored for efficient and collaborative use of denoising diffusion probabilistic models, CollaFuse enables shared server training and inference, alleviating client computational burdens. This is achieved by retaining data and computationally inexpensive GPU processes locally at each client while outsourcing the computationally expensive processes to the shared server. Demonstrated in a healthcare context, CollaFuse enhances privacy by highly reducing the need for sensitive information sharing. These capabilities hold the potential to impact various application areas, such as the design of edge computing solutions, healthcare research, or autonomous driving. In essence, our work advances distributed machine learning, shaping the future of collaborative GenAI networks.
category:
Short Paper
@@
Generative Models, Distributed Learning, Split Learning, Diffusion Model.
Domenique Zipperling, University of Bayreuth & Fraunhofer FIT
Simeon Allmendinger, University of Bayreuth & Fraunhofer FIT
Lukas Struppek, Technical University of Darmstadt & German Center for Artificial Intelligence
Niklas Kühl, University of Bayreuth & Fraunhofer FIT
COLLAFUSE: NAVIGATING LIMITED RESOURCES AND PRIVACY IN COLLABORATIVE GENERATIVE AI
{hangparas}
1em1
Chapter 0 Introduction
In the realm of Information Systems Research, the emergence of generative artificial intelligence (GenAI) technologies like ChatGPT and DALL-E has marked a significant milestone. These advancements, as detailed by [7, ], have broadened public access to GenAI and catalyzed its integration into diverse sectors. Among GenAI innovations, the denoising diffusion probabilistic model (DDPM) stands out for its ability to generate high-quality images through an advanced denoising process, outperforming earlier methods such as generative adversarial networks (GANs) [8] or variational autoencoders (VAEs) [16] in terms of diversity and convergence guarantees [21, 5]. However, the implementation of DDPMs in business analytics and other fields is not without challenges. These models demand extensive data sets and computational resources [30], which are often limited, especially in decentralized systems [12]. The healthcare sector exemplifies these constraints, where data scarcity, privacy concerns, and high costs of data collection are present [27, 29]. To address these challenges, researchers have explored various strategies, including patch-wise training [30], few-shot learning  [19, 23, 32], and notably, federated learning (FL) [20, 6]. FL enhances data accessibility, yet raises privacy concerns [24, 33] and the need for significant local computational capabilities.
In response, we introduce CollaFuse, a new collaborative learning and inference framework for DDPMs, inspired by split learning [9]. CollaFuse aims to balance the computationally intensive denoising process between local clients and a shared server, with a focus on optimizing the trade-off between performance, privacy, and resource utilization—which are crucial requirements for real-world information systems implementations. This framework transforms the optimization challenge into a multi-criteria problem, addressing the core requirements of such applications in practice. Building upon these criteria, our research investigates the impact of different degrees of collaboration defined by a cut-ratio $c\in[0,1]$, formulating two key hypotheses:
@@
Our initial analysis, including a healthcare-focused experiment with MRI brain scans, supports these hypotheses. As a consequence, CollaFuse holds promise for applicants such as small medical institutions or even individual practitioners with edge devices to engage in collaborative model training and inference, e.g., medical training. Beyond healthcare, the framework exhibits potential in domains like autonomous driving, where edge computing resources are constrained, yet computational demands and privacy considerations are high. Moving forward, our research will delve deeper into the CollaFuse framework, analyzing its performance in terms of image fidelity and diversity, assessing potential privacy risks, and exploring resource efficiency in additional scenarios. In pursuit of a more comprehensive understanding and depth our scenarios are characterized by a parameter grid: number of clients, data domain, absolute amount of client data, conditional inference, as well as client-dependent variance scheduler and cut-ratio.
This comprehensive investigation aims to advance our understanding of (distributive) GenAI within socio-technical systems as demanded by related literature [7, 1]. Consequently, our ongoing research on CollaFuse can offer guidelines on how to apply DDPMs collaboratively and work as a blueprint for future GenAI networks designs in various domains.
Chapter 1 Background
The foundations of our work are built by the collaborative concepts of distributed learning as well as the architectural principles of diffusion models. Federated Learning (FL) was first introduced by [20, ] as a solution to make distributed data accessible for training without storing it centrally. Since then, FL has been on a triumphant march and received a lot of attention [10, 15]. At its core, FL iteratively composes locally trained models into a global model requiring clients to share model updates and gradients, thereby increasing the privacy risk of, e.g., membership inference [24] or reconstruction attacks [33]. Furthermore, FL comes with high computational requirements at the client-side [26].
Another paradigm for collaborative learning is Split Learning (SL), which was introduced by [9, ] exploiting the sequencing of operations in neural networks. In general, SL splits a neural network among multiple clients and a shared server. This is especially intriguing as clients can use a server to train most of the model while keeping the data and labels locally.
@@
In 2020, research advanced diffusion models with the introduction of Denoising Diffusion Probabilistic Models (DDPM) [13], offering an alternative to GANs [8] for image generation. DDPMs involve two processes: diffusion and denoising. The diffusion process adds noise from a Gaussian distribution incrementally to an image over $T$ steps. In the denoising process, the model estimates the noise added at each step $t\in[0,T]$. While training the model, weights are updated, calculating the loss based on the difference between true and estimated noise. Accordingly, the initial training image is not necessary for the denoising process. On this basis, DDPMs are able to generate new images from pure Gaussian noise, which closely resemble the images of the training data set. Fidelity metrics, like the Kernel-Inception distance (KID) [4], gauge diffusion model performance by quantifying the difference between the distributions of real and generated images.
Chapter 2 Related Work
With Generative Adversarial Networks (GANs) being composed of two components, a generator and discriminator, research demonstrates different integration of the two components within the collaboration process. [11, ], being among the first to use distributed data sets to train GANs, propose MD-GAN. The approach consists of one central generator while clients hold discriminators locally. In contrast, [28, ] apply FL for GANs to cross-modality brain image synthesis. Using differential privacy gradient descent, only the generators of each client are aggregated. The discriminators remain local. [6, ] offer empirical results for FL-GANs indicating that synchronizing discriminator and generator across clients yields the best results for two data sets. On this basis, [18, ] offer IFL-GAN, an improved version of the FL-GAN, enabling local GANs to hold different weights resulting in faster convergence of the global model. As FL exhibits privacy concerns, there have been efforts to address these issues within collaborative training of GANs. [2, ] apply an algorithm for differentially private federated GANs to effectively tackle commonly occurring data issues. [27, ] address trust-related weaknesses of existing federated GAN solutions by combining three building blocks federated GANs, consortium blockchains, and Shamir’s Secret Sharing algorithm enabling generation of synthetic data in decentralized settings. Additionally, [31, ] propose a hybrid federated split learning framework for wireless networks and analyze the trade-off between training time and energy consumption, showing efforts to combine FL and SL.
Research on collaborative training methods is still scarce in the domain of Diffusion Models. [14, ] introduce the Phoenix technique for training unconditional diffusion models in a horizontal FL setting. The objective is to address mode coverage issues often seen in non-independent and identically distributed datasets. The data-sharing approach achieves a performance boost by sharing only 4-5% of the data among clients, minimizing communication overhead. Personalization and threshold filtering techniques outperform comparison methods in terms of precision and recall but fall short in image quality compared to the proposed technique. The paper suggests further exploration to enhance image quality in future work.
By mainly focusing on FL, current literature on collaborative GenAI, especially for DDPMs, neglects benefits from different collaborative paradigms. Adapting principles from SL can bring typical advantages such as reduced local resources and increased privacy to collaborative GenAI. In proposing CollaFuse we want to tap into these advantages and push GenAI to its next evolutionary level.
@@
Chapter 3 Framework
We propose CollaFuse, a framework facilitating collaborative GenAI for image generation with DDPMs across clients. Inspired by SL, the less resource-intensive diffusion process is computed locally by each client, whereas the computationally intensive denoising process is strategically split at step $t_{c}=(1-c)T$ during both training and inference. This results in a shared model (backbone) hosted on a shared server, coupled with individual local models for each client. The unique design of CollaFuse allows clients to retain potentially sensitive data locally while outsourcing the majority of computationally intensive denoising operations to a centralized server. The split is governed by the cut-ratio $c$, dictating the computational load for each client and the extent of disclosed information. For example, if $T=50$ and $c=0.8$, 20% of the denoising process (from $t=1$ to $t_{c}=10$) is trained and stored on the shared server, while the remaining 40 denoising steps (80%) are trained locally, ensuring privacy of these steps.
As illustrated in Figure 1a, from the client’s perspective, the training sequence orchestrated by CollaFuse comprises six key steps. The server initiates the client’s diffusion process (1), after which the client computes its image data’s diffusion process batch-wise (2) and forwards the resulting noised images, along with the corresponding added random noise $\epsilon=(\epsilon_{0},...,\epsilon_{t_{c}})$ for each image, back to the server (3). Utilizing the noised images, the server undertakes the initial phase of the denoising process, computing the loss between the estimated and provided random noise (4). The partially denoised images are subsequently transmitted back to the client (5), where the remaining denoising process is executed (6).
The multi-client perspective is depicted in Figure 1b, showcasing two clients providing noised images to the shared server and receiving partially denoised images for further training. Similarly, each client has the capability to generate new images by bypassing the diffusion process of existing images and solely denoising them from pure random noise.
Chapter 4 Experimental Evaluation
To assess our framework, CollaFuse, we simulate a healthcare-related scenario involving three clients and one server. Every client data set is independent comprising 4,920 MRI scans from 123 patients each. The hold-out test data set contains 5,000 images from 125 further patients [3]. The applied DDPM employs an identical cosine variance scheduler, $T=100$ steps, and maintains a fixed image size of ($128\times 128$) across clients. The training process spans $300$ epochs with a fixed learning rate of $0.001$ and a batch size of $150$.
Our experimental investigation delves into the impact of the cut-ratio $c$ on the trade-off between performance, disclosed information, and GPU energy consumption considering cut-ratio values $c\in[0.0,0.2,0.4,0.6,0.8,1.0]$. Performance is assessed using the common fidelity metric KID [4] on both the client-dependent training and hold-out data sets, employing the feature extractor from the clean-fid library [22]. GPU energy consumption is measured using the codecarbon Python package  [^1] https://mlco2.github.io/codecarbon/. Disclosed information for each client is approximated through the mean squared error (MSE) for a pixel-by-pixel comparison and KID scores between partially denoised images at the split step $t_{c}$ and real images of clients. To effectively simulate a distributed healthcare-related scenario, the experiment runs on a cluster with four NVIDIA A100-SXM4-40GB GPUs. Computation tasks are distributed across clients and the server, with each client utilizing a single GPU.
Chapter 5 Results
Our findings delve into the nuanced trade-off among all three dimensions: the performance, gauged by the fidelity of generated images; the extent of disclosed information, verifying whether images at the split point equal original images within respective client data sets; and the GPU power usage across clients and server. Figure 1 illustrates the trade-off between performance and disclosed information. Both are presented featuring stacked KID scores composed of all three clients. Additionally, disclosed information is assessed with the MSE metric. Lower KID scores signify enhanced performance, underscoring fidelity in generated images. Conversely, the investigation of disclosed information operates in the opposite direction. Here, the objective is to maximize KID and MSE scores, ensuring the revelation of as few characteristics as possible on the shared server to uphold stringent data privacy.
Figure 1: Experimental outcomes for performance and disclosed information in the CollaFuse framework based on our healthcare-related experiment: Figures include KID scores and MSE from pixel-wise comparisons. Collaborative efforts, illustrated by the U-shaped pattern of stacked KID scores, lead to better performances compared to exclusive local or global model training. Despite conducting up to 80% of the computationally intensive denoising steps on the server, a significant amount of information in the images remains concealed.
@@
The analysis of Figure 1 unveils two significant observations. Firstly, the stacked KID scores of performance exhibit a U-shaped pattern concerning the cut-ratio, particularly evident for client data. Consequently, collaborative efforts in our experiment lead to a reduced aggregated KID score compared to the scenario where clients exclusively train the model locally ($100\%$). Interestingly, the performance is reduced, if a large part of the global denoising process is conducted on the shared server ($0\%-40\%$), thereby contributing to the observed U-shaped trend. With regard to the disclosed information, both pixel-wise comparison and the KID score imply that, despite conducting up to 80% of computationally intensive denoising steps on the server, a substantial portion of information associated with the images remains concealed in comparison to total global denoising ($0\%$). Figure 1 exemplary illustrates the denoising process for different cut-ratios, which increases the comprehension of our results. Moreover, GPU power usage of the diffusion process exhibits limited computational intensity in the experiment, and the relocation of denoising steps to the server correlates positively with reduced local GPU energy demand.
Overall, our experiment supports Hypothesis 1, showing that collaborative learning with CollaFuse improves image fidelity compared to non-collaborative local training (c = 1). Hypothesis 2 is evident in disclosed information, and reduced local computational intensity when denoising steps are moved to the server. Fascinatingly, concerning performance, the hypothesis only holds to some extent as the results indicate a tipping point where further collaboration does not lead to an increase in performance.
Chapter 6 Conclusion and Outlook

### v2

@@ Abstract
In the landscape of generative artificial intelligence, diffusion models present challenges for socio-technical systems in data requirements and privacy. Traditional approaches like federated learning distribute the learning process but strain individual clients, especially with constrained resources. In response to these challenges, we introduce CollaFuse, a novel framework inspired by split learning. Tailored for efficient and collaborative use of denoising diffusion probabilistic models, CollaFuse enables shared server training and inference, alleviating client computational burdens. This is achieved by retaining data and computationally inexpensive GPU processes locally at each client while outsourcing the computationally expensive processes to the shared server. Demonstrated in a healthcare context, CollaFuse enhances privacy by reducing the need for sensitive information sharing. These capabilities hold the potential to impact various application areas, such as edge computing, healthcare research, or autonomous driving. In essence, our work advances distributed machine learning, shaping the future of collaborative GenAI networks.
@@
0 Introduction
In the realm of Information Systems Research, the emergence of generative artificial intelligence (GenAI) technologies like ChatGPT and DALL-E has marked a significant milestone. These advancements, as detailed by feuerriegel2023generative, have broadened public access to GenAI and catalyzed its integration into diverse sectors. Among GenAI innovations, the denoising diffusion probabilistic model (DDPM) stands out for its ability to generate high-quality images through an advanced denoising process, outperforming earlier methods such as generative adversarial networks (GANs) (goodfellow20gans) or variational autoencoders (VAEs) (Kingma2014VAE) in terms of diversity and convergence guarantees (nichol21improving; dhariwal21beating). However, the implementation of DDPMs in business analytics and other fields is not without challenges. These models demand extensive data sets and computational resources (wang2023patch), which are often limited, especially in decentralized systems (hirt2019cognition). The healthcare sector exemplifies these constraints, where data scarcity, privacy concerns, and high costs of data collection are present (Veeraragavan2023SecuringFLGAN; Wang2023PrivacyFL_ICU). Nonetheless, DDPMs possess significant potential in the healthcare sector for example in medical image synthesis, reconstruction, or Image-to-Image translation. (KAZEROUNI2023102846). To address these challenges, researchers have explored various strategies, including patch-wise training (wang2023patch), few-shot learning  (Lu2023specialist; ruiz2023dreambooth; Zhang2023Sine), and notably, federated learning (FL) (McMahan2017FL; fan2020federated). FL enhances data accessibility, yet raises privacy concerns (shokri17meminf; zhu19leakage) and the need for significant local computational capabilities.
In response, we introduce CollaFuse, a new collaborative learning and inference framework for DDPMs, inspired by split learning (SL) (gupta2018distributed). CollaFuse aims to balance the computationally intensive denoising process between local clients and a shared server, with a focus on optimizing the trade-off between performance, privacy, and resource utilization—which are crucial requirements for real-world information systems implementations. This framework transforms the optimization challenge into a multi-criteria problem, addressing the core requirements of such applications in practice. Building upon these criteria, our research investigates the impact of different degrees of collaboration defined by a cut-ratio $c\in[0,1]$, formulating two key hypotheses:
@@ 0 Introduction
Our initial analysis, including a healthcare-focused experiment with magnetic resonance imaging (MRI) brain scans, supports these hypotheses. As a consequence, CollaFuse holds promise for applicants such as small medical institutions or even individual practitioners with edge devices to engage in collaborative model training and inference, e.g., medical training. Beyond healthcare, the framework exhibits potential in domains like autonomous driving, where edge computing resources are constrained, yet computational demands and privacy considerations are high. Moving forward, our research will delve deeper into the CollaFuse framework, analyzing its performance in terms of image fidelity and diversity, assessing potential privacy risks, and exploring resource efficiency in additional scenarios. This comprehensive investigation aims to advance our understanding of (distributive) GenAI within socio-technical systems as demanded by related literature (feuerriegel2023generative; abbasi2023data). Consequently, our ongoing research on CollaFuse can offer guidelines on how to apply DDPMs collaboratively and work as a blueprint for future GenAI networks designs in various domains.
1 Background and Related Work
@@
Our research is grounded in the collaborative concepts of distributed learning and the architectural principles of diffusion models. FL was first introduced by McMahan2017FL utilizing distributed data for training without storing it centrally. Since then, FL has been on a triumphant march and received a lot of attention (hard2019federated). At its core, FL iteratively composes locally trained models into a global model requiring clients to share model updates and gradients, which increases the risk of data leakage (shokri17meminf; zhu19leakage). Furthermore, FL comes with high computational requirements at the client-side (SplitFed2022). Another paradigm for collaborative learning is SL exploiting the sequencing of operations in neural networks (gupta2018distributed). In general, SL splits a neural network among multiple clients and a shared server. This is especially intriguing as clients can use a server to train most of the model while keeping the data and labels locally. In 2020, research advanced diffusion models with the introduction of Denoising Diffusion Probabilistic Models (DDPM) (Ho2020), offering an alternative to GANs (goodfellow20gans) for image generation. DDPMs involve two processes: diffusion and denoising. The diffusion process adds noise from a Gaussian distribution incrementally to an image over $T$ steps. In the denoising process, the model estimates the noise added at each step $t\in[0,T]$. While training the model, weights are updated, calculating the loss based on the difference between true and estimated noise. Accordingly, the initial training image is not necessary for the denoising process. On this basis, DDPMs are able to generate new images from pure Gaussian noise, which closely resemble the images of the training data set. Fidelity metrics, like the Kernel-Inception distance (KID) (binkowsi18kid), gauge diffusion model performance by quantifying the difference between the distributions of real and generated images.
Recent developments in GenAI have spurred interest within the information systems (IS) community.  feuerriegel2023generative and Banh2023GenAI broadly define GenAI, covering its architectures and applications, while identifying challenges and research questions for the IS community. Additionally, research has addressed how GenAI, particularly the language generation model ChatGPT, influences education marketing or intellectual property (Peres2023; Burger2023). Moreover, FL has also captured the attention of the IS community. karnebogen2023SLR_FL have introduced a taxonomy for FL applications, while Wang2023PrivacyFL_ICU have investigated FL for privacy-preserving healthcare applications. hirt2023enabling explore the impacts of distributed learning on organizations. However, the integration of FL with GenAI, particularly in image synthesis, remains unexplored within IS research, prompting a look towards approaches from other communities, such as computer science. Within the computer science community GANs have been used for a long time to synthesize images. As GANs are composed of two components, a generator and discriminator, research has been centered around whether both or only one component should be trained collaboratively. Hardy2019MD-GAN; wang2023fedmedgan propose only to train the generator collaboratively and to train the discriminator at each client. In contrast, fan2020federated offer empirical results indicating that synchronizing discriminator and generator across clients yields the best results, leading to the proposal of an improved FL-Gan version (Li2022IFL-GAN). As FL exhibits privacy concerns, augenstein2020generative apply differential privacy while Veeraragavan2023SecuringFLGAN add consortium blockchains and Shamir’s Secret Sharing algorithm to federated GANs. Additionally, Benshun2023SL-GAN propose a federated split learning framework showing efforts to combine FL and SL. Research on collaborative training methods is still scarce in the domain of Diffusion Models. jothiraj2023phoenix introduce Phoenix, a method for training diffusion models in federated learning to address data diversity issues, sharing minimal data (4-5%) among clients to cut communication costs. It excels in precision and recall but needs further development to enhance image quality. Current research on collaborative GenAI largely overlooks the advantages of other DDPMs and SL, which can offer reduced resource use and enhanced privacy. With CollaFuse, we aim to leverage these benefits to advance GenAI.
2 Framework
We propose CollaFuse, a framework facilitating collaborative GenAI for image generation with DDPMs across clients. Inspired by SL, the less resource-intensive diffusion process is computed locally by each client, whereas the computationally intensive denoising process is strategically split at step $t_{c}=(1-c)T$ during both training and inference. This results in a shared model (backbone) hosted on a shared server, coupled with individual local models for each client. The unique design of CollaFuse allows clients to retain sensitive data locally while outsourcing the majority of computationally intensive denoising operations to a centralized server. The split is governed by the cut-ratio $c$, dictating the computational load for each client and the extent of disclosed information. Figure 1 shows this governance, where, for instance, with $T=100$ and $c=0.8$, the first 20% of the denoising process (from $t=1$ to $t_{0.8}=20$) is managed on a shared server and the subsequent 80 steps (80%) are locally executed, maintaining the privacy of these stages
As illustrated in Figure 1a, from the client’s perspective, the training sequence orchestrated by CollaFuse comprises six key steps. The server initiates the client’s diffusion process (1), after which the client computes its image data’s diffusion process batch-wise (2) and forwards the resulting noised images, along with the corresponding added random noise $\epsilon=(\epsilon_{0},...,\epsilon_{t_{c}})$ for each image, back to the server (3). Utilizing the noised images, the server undertakes the initial phase of the denoising process, computing the loss between the estimated and provided random noise (4). The partially denoised images are subsequently transmitted back to the client (5), where the remaining denoising process is executed (6). The system perspective is depicted in Figure 1b, showcasing two clients providing noised images to the shared server and receiving partially denoised images for further training. Further, clients can generate new images by sampling images from pure random noise and passing them to the server.
3 Experimental Evaluation
To assess our framework, CollaFuse, we simulate a healthcare-related scenario involving three clients and one server. Every client data set is independent comprising 4,920 MRI scans from 123 patients each. The hold-out test data set contains 5,000 images from 125 further patients (Bakas2017). Our diffusion model employs a U-Net architecture (ronneberger15unet) including ResNet (resnet_he) blocks for down- and upsampling and self-attention mechanisms (vaswani17attention) for feature refinement and textual guidance. Furthermore, we employ an identical cosine variance scheduler, $T=100$ steps, and maintain a fixed image size of ($128\times 128$) across clients. The training process spans $300$ epochs with a fixed learning rate of $0.001$ and a batch size of $150$. Our investigation delves into the impact of the cut-ratio $c$ on the trade-off between performance, disclosed information, and GPU energy consumption considering cut-ratio values $c\in[0.0,0.2,...,1.0]$. Performance is assessed using the common fidelity metric KID (binkowsi18kid) on both the client-dependent training and hold-out data sets, employing the feature extractor from the clean-fid library (parmar_On). GPU energy consumption is measured using the codecarbon Python package  [^1] https://mlco2.github.io/codecarbon/. Disclosed information for each client is approximated through the mean squared error (MSE) for a pixel-by-pixel comparison and KID scores between partially denoised images at the split step $t_{c}$ and real images of clients. To simulate a distributed healthcare-related scenario, we utilize a cluster of four NVIDIA A100-SXM4-40GB GPUs. The implementation of the architecture, framework and experiment can be found here: https://github.com/SimeonAllmendinger/collafuse.git.
@@ 3 Experimental Evaluation
4 Results
@@
Our findings analyze trade-off among all three dimensions: the performance, gauged by the fidelity of generated images; the extent of disclosed information, verifying whether images at the split point equal original images within respective client data sets; and the GPU power usage across clients and server. Figure 1 illustrates the trade-off between performance and disclosed information. We calculate the KID score to assess generative performance and out-of-sample robustness (upper-left, lower-left). Information disclosure is evaluated by comparing original client data to server-generated images, using KID for distribution analysis (upper-right) and MSE for pixel-level comparison. Lower KID scores signify enhanced performance, underscoring fidelity in generated images. Conversely, for disclosed information the objective is to maximize KID and MSE scores, ensuring the revelation of as few characteristics as possible on the shared server. The analysis of Figure 1 unveils two significant observations. Firstly, the stacked KID scores of performance exhibit a U-shaped pattern concerning the cut-ratio, particularly evident for client data. Consequently, collaborative efforts may lead to a reduced aggregated KID score compared to training the models locally ($100\%$). Interestingly, the performance is reduced, if a large part of the global denoising process is conducted on the shared server ($0\%-40\%$), thereby contributing to the observed U-shaped trend. With regard to the disclosed information, both pixel-wise comparison and the KID score imply that, despite conducting up to 80% of computationally intensive denoising steps on the server, a substantial portion of information associated with the images remains concealed in comparison to total global denoising ($0\%$). Moreover, GPU power usage of the diffusion process exhibits limited computational intensity in the experiment, and the relocation of denoising steps to the server correlates positively with reduced local GPU energy demand. Overall, our experiment supports Hypothesis 1, showing that collaborative learning with CollaFuse improves image fidelity compared to non-collaborative local training (c = 1). Hypothesis 2 is evident in disclosed information, and reduced local computational intensity when denoising steps are moved to the server. Fascinatingly, concerning performance, the hypothesis only holds to some extent as the results indicate a tipping point where further collaboration does not lead to an increase in performance.
Figure 1: Preliminary results: We calculate the KID score to assess generative performance and out-of-sample robustness (upper-left, lower-left). Information disclosure is evaluated by comparing original client data to server-generated images, using KID for distribution (upper-right) and MSE for pixel-level comparison. Despite conducting up to 80% of the computationally intensive denoising steps on the server, a significant amount of information in the images remains concealed.
5 Conclusion and Outlook